# Origins of the UV continuum and Balmer emission lines in Little Red Dots: observational validation of dense gas envelope models enshrouding the AGN

### v1

@@ I Introduction
A newly discovered population of broad Balmer line emitters known as little red dots (LRDs; Matthee et al. 2024), presumably involving active galactic nuclei (AGNs), is one of the most intriguing findings from the James Webb Space Telescope (JWST). LRDs predominantly appear at high redshifts () and rapidly decline in abundance toward lower redshifts (e.g., Kocevski et al., 2025; Ma et al., 2025; Zhuang et al., 2025). Their unique spectral and morphological properties set them apart from both typical low-redshift AGNs and luminous quasars (e.g., Greene et al., 2024; Labbé et al., 2024; Furtak et al., 2024b; Akins et al., 2025; Hviding et al., 2025; Maiolino et al., 2025; Ji et al., 2025). LRDs may represent a population of black holes (BHs) with masses , caught in a transient phase of rapid growth from seed BHs to supermassive regimes (Inayoshi, 2025).
@@
The presence of characteristic Balmer-transition signatures in LRDs, including prominent absorption superimposed on broad Balmer emission lines, unusually large Balmer decrements, deep Balmer breaks, large equivalent widths of broad H emission, and exponential line profiles of the broad line components suggest that LRDs are AGNs embedded in dense gas, more specifically broad line regions (BLRs) residing within dense gaseous structures (e.g., Juodžbalis et al., 2024; Lin et al., 2024; Inayoshi and Maiolino, 2025; Ji et al., 2025; de Graaff et al., 2025c; Naidu et al., 2025; Rusakov et al., 2025b). When such dense gas clumps in BLRs become optically thick and act as Balmer line and continuum emitters, their thermal emission yields an optically red continuum spectrum by the Wien tail of a blackbody spectrum with an effective temperature of , without dust reddening (Inayoshi et al., 2025b; Kido et al., 2025; Liu et al., 2025; Lin et al., 2025; Begelman and Dexter, 2025). This interpretation is novel and critical, as JWST/MIRI photometry disfavors significant rest-frame near- to mid-infrared emission from AGN-heated hot dust (Akins et al., 2025; Casey et al., 2025; Chen et al., 2025b). As additional consequences, dense nuclear gas can efficiently cool X-ray emitting hot coronae via Compton scattering (Yue et al., 2024; Maiolino et al., 2025) and dissipates the power of nascent jets within the gaseous envelope (Mazzolari et al., 2024; Gloudemans et al., 2025).
@@
This AGN+dense gas envelope scenario, where the red optical continuum originates from gaseous thermal emission or from selective absorption of Balmer limit photons, successfully explains many of the puzzling properties of LRDs. However, it faces an important question regarding the origin of the observed UV continuum and broad-line emissions (see also Inayoshi and Ho, 2025). If the gas envelope fully obscures the UV radiation emitted from the accretion disk, the UV continuum of LRDs should arise from external components, such as compact young stellar populations. In this case, depending on the gas clumpiness and the optical depths to line transitions, the BLRs can be powered either by ionizing radiation from the the stellar population (Model A; Inayoshi et al., 2025a) or by radiation reprocessed from the embedded AGN that leaks out after multiple scattering (Model B; Naidu et al., 2025). Model (A) requires relatively intense star formation to supply sufficient ionizing photons to sustain the BLRs, whereas Model (B) demands a specific density and clumpiness structure in which the envelope is opaque to the UV continuum but allows Balmer lines to escape. Alternatively, if UV radiation produced from the accretion disk can escape preferentially toward the polar directions owing to rotation or anisotropic gas geometry, these photons could directly photoionize the BLRs as in the standard AGN picture (Model C; Lin et al., 2025).
@@ II Data and sample selection
We utilize all JWST/NIRSpec Prism spectra available on the DAWN JWST Archive (DJA; Brammer and Valentino, 2025). Spectra on the DJA version 4.4 are reduced with msaexp (Brammer, 2023), following de Graaff et al. (2025a) and Heintz et al. (2024), and the 1D spectra are extracted using the optimal aperture method (Horne, 1986). We only used grade=3 NIRSpec/Prism spectra without missing data due to the detector gap from rest far-UV to optical bands. We focus on the NIRSpec/Prism sample, because one of our goals is to spectroscopically measure the rest UV continuum colors and luminosity. Requiring NIRSpec grating observations would enable us to capture more fainter and/or narrower broad line components, but it would make our sample much more smaller and heterogeneous. This selection includes 80,367 spectra in total.
@@
We next select LRDs among the broad line emitters based on their colors and compactness. Following Hviding et al. (2025), we define the V-shaped LRD colors as:
@@
where and are rest UV and optical slopes measured from the spectra at and , respectively, with masking out wavelength ranges where strong emission lines are expected. Figure 2 (a) presents the distribution of the broad H emitters at , and LRDs are marked by red in this diagram. We finally select compact sources satisfying . We obtain a sample of 28 LRDs in total, including observations from CANUCS (Sarrouh et al., 2025), CAPERS (GO-6368; PI: Dickinson), CEERS (Finkelstein et al., 2025), JADES (Curtis-Lake et al., 2025; Scholtz et al., 2025), NEXUS (Shen et al., 2024), RUBIES (de Graaff et al., 2025a), UNCOVER (Bezanson et al., 2024), GTO-Wide (Maseda et al., 2024), and other small programs (GO-2198; PI: Barrufet and DD-6585; PI: Coulter). Figure 2 (b) shows an example of LRDs.
@@
and the same compactness criterion, among the broad H emitters (Figure 2 (a) blue). Hereafter, we refer to this control sample as Little Blue Dots (LBDs), and use it to probe the effect of enshrouding envelopes in LRDs. We obtain a sample of 9 LBDs, including observations by Abdurro’uf et al. (2024) and Wang et al. (2024) in addition to those for LRDs. An LBD example is shown in Figure 2 (c).
For all LRDs and LBDs, we measure the UV luminosity and broad+narrow H line luminosity from the NIRSpec/Prism spectrum. We first fit the continuum around the line with a power-law function, and fit a two-component Gaussian profile on the observed H line profile subsequently. The two-component model has five free parameters (the redshift and line width and line luminosity for both narrow+broad components). We use the python implementation of MCMC (emcee; Foreman-Mackey et al., 2013) to estimate the posteriors of the five free parameters, taking the 50th percentile as the best estimation and 16th- to 84th-percentile range as the uncertainties. The UV luminosity is directly measured from the spectrum at Å.
We also apply a dust attenuation correction, so that we can examine the intrinsic energy budget between rest UV light and ionizing radiation and explore the ionizing sources in LRDs and LBDs. Although their UV colors are relatively blue by definition, they are not as blue as dust-free () and suggest a modest dust attenuation. We therefore use the rest UV color as the probe of dust attenuation, assuming the Calzetti extinction law (; Calzetti et al., 2000). The dust attenuation estimated from the UV slope is typically small for LRDs ( mag), which does not violate the non-detection of hot/cold dust emissions of LRDs in JWST/MIRI or ALMA (Casey et al., 2025; Chen et al., 2025b).
@@ III.1 Energy balance between UV and broad/narrow lines in LRDs and LBDs
Figure 3 presents the UV and H luminosities of LRDs (top) and LBDs (bottom) with the broad and narrow H components separately in the left and right panels. For reference, we overlay the UV-to-H luminosity ratios for young starburst galaxies (black dotted) and for local AGNs (solid red). The UV-to-H luminosity ratio for young starburst galaxies is computed based on BPASS stellar population models (Eldridge et al., 2017) combined with photoionization calculations performed with Cloudy (v23.01; Chatzikos et al., 2023). Specifically, we compute the ratios by varying the electron density and stellar age over the ranges of and yr. For AGNs, the UV-to-H luminosity ratios are measured directly from the compiled Type 1 AGN spectrum (Vanden Berk et al., 2001). We decompose the H line profile into narrow and broad components by fitting two Gaussian profiles, and derive the corresponding ratios for each component separately. Table 1 collates observations and predictions.
@@ III.2 Lyman-alpha emission in LRDs and LBDs
We also examine the Ly emission lines in the sample broad line emitters. Since Ly is a resonance line, Ly photons are unlikely to escape through a dense gas envelope. Therefore, the Ly occurrence rate and the line strength provide a useful diagnostic of the origin of (narrow) hydrogen recombination lines. We use the best-fit power-law for the UV continuum (obtained in Section II) as the continuum baseline around the Ly emission, and we model the Ly line with a single Gaussian on top of the continuum. The IGM attenuation is applied on the continuum following the prescription by Inoue et al. (2014) and Asada et al. (2025), so that the potential rounded shape of the Lyman break can also be parameterized. We then identify Ly emitters (LAEs) among the sample broad H emitters by searching for Ly lines with S/N. Sources with Ly detections are marked with squares in Figure 2 (a), and the two examples shown in panel (b) and (c) are both LAEs.
The Ly occurrence rates are (11/28) and (2/9) among LRDs and LBDs. These rates are statistically consistent with each other, or LRDs may have even higher occurrence rate than LBDs, indicating that the presence of a dense gas envelope does not suppress the observability of Ly photons in these two classes. Moreover, the occurrence rate of is comparable to that observed in normal star-forming galaxies (SFGs) at similar redshifts and UV magnitudes ( mag; e.g., Stark et al. 2011; De Barros et al. 2017; Kusakabe et al. 2020; Goovaerts et al. 2023; Napolitano et al. 2025). The similar Ly occurrence rates in LRDs, LBDs, and normal SFGs further suggest that the narrow lines originate outside the dense gas envelope, most plausibly from H ii regions within the host galaxies.
Figure 4 presents the Ly equivalent width of the sample, plotted against their rest UV absolute magnitudes. No systematic difference can be seen between LRDs and LBDs, and they occupy very similar locus in the EW(Ly)- plot as normal SFGs at the same radshift range identified with VLT/MUSE (Kerutt et al., 2022). Considering AGN-powered Ly emission can be more luminous than normal SFGs (e.g., Sobral et al., 2018), there seems no strong evidence supporting that the (narrow) Ly lines in LRDs or LBDs are mainly powered by AGN accretion. We should caution that, however, our search for the Ly emission with NIRSpec/Prism spectra could be incomplete particularly for low EW(Ly) sources due to the low spectral resolution of the Prism at these short wavelengths. It is also difficult to compare the intrinsic emission line profiles between Ly and H narrow component due to the low spectral resolution, which could otherwise be informative to probe the origin of the Ly line engine and escape mechanisms (Baek and Ferrara, 2013; Yang et al., 2014). Further large spectroscopic campaigns of LRDs and LBDs at rest UV wavelengths with higher spectral resolution are needed to improve the statistics.
@@ IV.1 Constraint on LRD models from spectroscopic observations
In this section, we discuss the constraint on the models for LRDs from the observational results presented in the previous section. The observed correlations between the UV continuum and H lines, together with their luminosity ratios and Ly occurrence rates, provide key diagnostics of the dominant ionizing sources. We examine whether various envelope models in the literature can consistently account for these observational constraints and assess their relative plausibility. We mainly focus on the envelope models proposed in literature (see Figure 1), but also briefly mention the possibility of other models than envelope models.
@@
(e.g., Kido et al., 2025; Begelman and Dexter, 2025).
@@
We also note that the stellar origin for the UV continuum can account for observations of LRDs spatially resolved in the rest-frame UV, comprising of the samples reported in the literature (e.g., Rinaldi et al., 2025; Kokorev et al., 2025). The remaining LRDs are spatially unresolved, including gravitationally lensed sources with tighter size limits () yet show V-shaped SEDs (e.g., Furtak et al., 2024a; Chen et al., 2025a), which suggests that the UV emission mainly originates from young compact stellar clusters in the nuclear region, as described in this model (Figure 1a). Overall, this model provides a self-consistent explanation for the observational results and is thus the most favored among the envelope models considered here.
@@
(Figure 1b; Naidu et al. 2025; de Graaff et al. 2025c).
This scenario can explain the tight correlation of the UV and the narrow H luminosity in LRDs. However, it does not naturally account for the observed correlation between the UV continuum and the broad H luminosity.
@@
predominantly powered by the AGN accretion disk (e.g., Lin et al., 2025)111The non-spherical envelope model is defined as the case where both the UV continuum and the reprocessed emission lines powered by the AGN can escape through low-density, polar directions. Although Lin et al. (2025) adopt a non-spherical gas geometry and provide a schematic illustration, we here note that they do not explicitly specify the origin of the UV continuum, treating it as an open question..
@@
A possible explanation for the observed enhancement of H-to-UV luminosity ratio under this hypothesis is a higher gas covering fraction in the nuclear region during the LRD phase. Near the surface of the dense gaseous envelope (on spatial scales comparable to the BLR sizes inferred from reverberation mapping in nearby AGNs), BLR clouds occupy a large fraction of the solid angle of the radiation-escape funnel, thereby reprocessing ionizing photons into line emission more efficiently (e.g., Maiolino et al., 2025; Yan et al., 2025). Therefore, the non-spherical envelope model could explain the different UV-to-H luminosity ratios between LRDs and LBDs if the BLRs during the LRD phase are assumed to have higher covering fractions and larger gas column densities than normal unobscured AGNs (LBDs). However, this assumption introduces a larger scatter in the UV-H relation and requires LBDs to have lower covering fractions. It also remains physically unclear why the offset between LRDs and LBDs appears at a factor of .
In any case, a non-spherical envelope phase may still play an important role during the evolutionary transition from LRDs to normal AGNs (see e.g., Mérida et al., 2025). The continuous distribution observed in the color–color diagram (Figure 2a) from LRDs to LBDs suggests that LRDs can gradually evolve into normal Type 1 AGNs. Such a transition could naturally proceed through a non-spherical envelope phase, providing a physical link to the unified model of AGNs (e.g., Netzer, 2015).
@@
Although recent theoretical and observational works often focus on envelope models, they were originally proposed to explain a subset of LRDs with extreme properties such as deep Balmer breaks, Balmer absorption lines, exponential wing line profiles, or blackbody-like continuum shape from rest optical to infrared (e.g., Naidu et al., 2025; Rusakov et al., 2025b; Lin et al., 2025). It is thus possible that the dense gas envelope is needed only for these extreme LRDs and different physics may lie behind the overall normal population of LRDs.

### v2

@@ I Introduction
A newly discovered population of broad Balmer line emitters known as little red dots (LRDs; J. Matthee et al. 2024), presumably involving active galactic nuclei (AGNs), is one of the most intriguing findings from the James Webb Space Telescope (JWST). LRDs predominantly appear at high redshifts () and rapidly decline in abundance toward lower redshifts (e.g., D. D. Kocevski et al., 2025; Y. Ma et al., 2025; M.-Y. Zhuang et al., 2025). Their unique spectral and morphological properties set them apart from both typical low-redshift AGNs and luminous quasars (e.g., J. E. Greene et al., 2024; I. Labbé et al., 2024; L. J. Furtak et al., 2024a; H. B. Akins et al., 2025; R. E. Hviding et al., 2025; R. Maiolino et al., 2025; X. Ji et al., 2025). LRDs may represent a population of black holes (BHs) with masses , caught in a transient phase of rapid growth from seed BHs to supermassive regimes (K. Inayoshi, 2025).
@@
The presence of characteristic Balmer-transition signatures in LRDs, including prominent absorption superimposed on broad Balmer emission lines, unusually large Balmer decrements, deep Balmer breaks, large equivalent widths of broad H emission, and exponential line profiles of the broad line components suggest that LRDs are AGNs embedded in dense gas, more specifically broad line regions (BLRs) residing within dense gaseous structures (e.g., I. Juodžbalis et al., 2024; X. Lin et al., 2024; K. Inayoshi & R. Maiolino, 2025; X. Ji et al., 2025; A. de Graaff et al., 2025a; R. P. Naidu et al., 2025; V. Rusakov et al., 2025a). When such dense gas clumps in BLRs become optically thick and act as Balmer line and continuum emitters, their thermal emission yields an optically red continuum spectrum by the Wien tail of a blackbody spectrum with an effective temperature of , without dust reddening (K. Inayoshi et al., 2025b; D. Kido et al., 2025; H. Liu et al., 2025; X. Lin et al., 2025; M. C. Begelman & J. Dexter, 2025). This interpretation is novel and critical, as JWST/MIRI photometry disfavors significant rest-frame near- to mid-infrared emission from AGN-heated hot dust (H. B. Akins et al., 2025; C. M. Casey et al., 2025; K. Chen et al., 2025). As additional consequences, dense nuclear gas can efficiently cool X-ray emitting hot coronae via Compton scattering (M. Yue et al., 2024; R. Maiolino et al., 2025) and dissipates the power of nascent jets within the gaseous envelope (G. Mazzolari et al., 2024; A. J. Gloudemans et al., 2025).
@@
This AGN+dense gas envelope scenario, as known as “BH* model” in literature, is originally proposed by R. P. Naidu et al. (2025) to explain the red optical continuum with gaseous thermal emission or with selective absorption of Balmer limit photons, and it successfully explains many of the puzzling properties of LRDs. However, it faces an important question regarding the origin of the observed UV continuum and broad-line emissions (see also K. Inayoshi & L. C. Ho, 2025), and different types of “BH* model” have been discussed. If the gas envelope fully obscures the UV radiation emitted from the accretion disk, the UV continuum of LRDs should arise from external components, such as compact young stellar populations. In this case, depending on the gas clumpiness and the optical depths to line transitions, the BLRs can be powered either by ionizing radiation from the the stellar population (Model A; K. Inayoshi et al., 2025a) or by radiation reprocessed from the embedded AGN that leaks out after multiple scattering (Model B; R. P. Naidu et al., 2025). Model (A) requires relatively intense star formation to supply sufficient ionizing photons to sustain the BLRs, whereas Model (B) demands a specific density and clumpiness structure in which the envelope is opaque to the UV continuum but allows Balmer lines to escape. Alternatively, if UV radiation produced from the accretion disk can escape preferentially toward the polar directions owing to rotation or anisotropic gas geometry, these photons could directly photoionize the BLRs as in the standard AGN picture (Model C; X. Lin et al., 2025).
@@ II Data and sample selection
We utilize all JWST/NIRSpec Prism spectra available on the DAWN JWST Archive (DJA; G. Brammer & F. Valentino, 2025). Spectra on the DJA version 4.4 are reduced with msaexp (G. Brammer, 2023), following A. de Graaff et al. (2025b) and K. E. Heintz et al. (2024), and the 1D spectra are extracted using the optimal aperture method (K. Horne, 1986). We only used grade=3 NIRSpec/Prism spectra without missing data due to the detector gap from rest far-UV to optical bands. We focus on the NIRSpec/Prism sample, because one of our goals is to spectroscopically measure the rest UV continuum colors and luminosity. Requiring NIRSpec grating observations would enable us to capture more fainter and/or narrower broad line components, but it would make our sample much more smaller and heterogeneous. This selection includes 80,367 spectra in total.
@@
We next select LRDs among the broad line emitters based on their colors and compactness. Following R. E. Hviding et al. (2025), we define the V-shaped LRD colors as:
@@
where and are rest UV and optical slopes measured from the spectra at and , respectively, with masking out wavelength ranges where strong emission lines are expected. Figure 2 (a) presents the distribution of the broad H emitters at , and LRDs are marked by red in this diagram. We finally select compact sources satisfying . We obtain a sample of 28 LRDs in total, including observations from CANUCS (G. T. E. Sarrouh et al., 2025), CAPERS (GO-6368; PI: Dickinson), CEERS (S. L. Finkelstein et al., 2025), JADES (E. Curtis-Lake et al., 2025; J. Scholtz et al., 2025), NEXUS (Y. Shen et al., 2024), RUBIES (A. de Graaff et al., 2025b), UNCOVER (R. Bezanson et al., 2024), GTO-Wide (M. V. Maseda et al., 2024), and other small programs (GO-2198; PI: Barrufet and DD-6585; PI: Coulter). Figure 2 (b) shows an example of LRDs.
@@
and the same compactness criterion, among the broad H emitters (Figure 2 (a) blue). Hereafter, we refer to this control sample as Little Blue Dots (LBDs), and use it to probe the effect of enshrouding envelopes in LRDs. We obtain a sample of 9 LBDs, including observations by Abdurro’uf et al. (2024) and F. Wang et al. (2024) in addition to those for LRDs. An LBD example is shown in Figure 2 (c).
For all LRDs and LBDs, we measure the UV luminosity and broad+narrow H line luminosity from the NIRSpec/Prism spectrum. We first fit the continuum around the line with a power-law function, and fit a two-component Gaussian profile on the observed H line profile subsequently. The two-component model has five free parameters (the redshift and line width and line luminosity for both narrow+broad components). We use the python implementation of MCMC (emcee; D. Foreman-Mackey et al., 2013) to estimate the posteriors of the five free parameters, taking the 50th percentile as the best estimation and 16th- to 84th-percentile range as the uncertainties. The UV luminosity is directly measured from the spectrum at Å.
We also apply a dust attenuation correction, so that we can examine the intrinsic energy budget between rest UV light and ionizing radiation and explore the ionizing sources in LRDs and LBDs. Although their UV colors are relatively blue by definition, they are not as blue as dust-free () and suggest a modest dust attenuation. We therefore use the rest UV color as the probe of dust attenuation, assuming the Calzetti extinction law (; D. Calzetti et al., 2000). The dust attenuation estimated from the UV slope is typically small for LRDs ( mag), which does not violate the non-detection of hot/cold dust emissions of LRDs in JWST/MIRI or ALMA (C. M. Casey et al., 2025; K. Chen et al., 2025).
@@ III.1 Energy balance between UV and broad/narrow lines in LRDs and LBDs
Figure 3 presents the UV and H luminosities of LRDs (top) and LBDs (bottom) with the broad and narrow H components separately in the left and right panels. For reference, we overlay the UV-to-H luminosity ratios for young starburst galaxies (black dotted) and for local AGNs (solid red). The UV-to-H luminosity ratio for young starburst galaxies is computed based on BPASS stellar population models (J. J. Eldridge et al., 2017) combined with photoionization calculations performed with Cloudy (v23.01; M. Chatzikos et al., 2023). Specifically, we compute the ratios by varying the electron density and stellar age over the ranges of and yr. For AGNs, the UV-to-H luminosity ratios are measured directly from the compiled Type 1 AGN spectrum (D. E. Vanden Berk et al., 2001). We decompose the H line profile into narrow and broad components by fitting two Gaussian profiles, and derive the corresponding ratios for each component separately. Table 1 collates observations and predictions.
@@ III.2 Lyman-alpha emission in LRDs and LBDs
We also examine the Ly emission lines in the sample broad line emitters. Since Ly is a resonance line, Ly photons are unlikely to escape through a dense gas envelope. Therefore, the Ly occurrence rate and the line strength provide a useful diagnostic of the origin of (narrow) hydrogen recombination lines. We use the best-fit power-law for the UV continuum (obtained in Section II) as the continuum baseline around the Ly emission, and we model the Ly line with a single Gaussian on top of the continuum. The IGM attenuation is applied on the continuum following the prescription by A. K. Inoue et al. (2014) and Y. Asada et al. (2025), so that the potential rounded shape of the Lyman break can also be parameterized. We then identify Ly emitters (LAEs) among the sample broad H emitters by searching for Ly lines with S/N. Sources with Ly detections are marked with squares in Figure 2 (a), and the two examples shown in panel (b) and (c) are both LAEs.
The Ly occurrence rates are (11/28) and (2/9) among LRDs and LBDs. These rates are statistically consistent with each other, or LRDs may have even higher occurrence rate than LBDs, indicating that the presence of a dense gas envelope does not suppress the observability of Ly photons in these two classes. Moreover, the occurrence rate of is comparable to that observed in normal star-forming galaxies (SFGs) at similar redshifts and UV magnitudes ( mag; e.g., D. P. Stark et al. 2011; S. De Barros et al. 2017; H. Kusakabe et al. 2020; I. Goovaerts et al. 2023; L. Napolitano et al. 2025). The similar Ly occurrence rates in LRDs, LBDs, and normal SFGs further suggest that the narrow lines originate outside the dense gas envelope, most plausibly from H ii regions within the host galaxies.
Figure 4 presents the Ly equivalent width of the sample, plotted against their rest UV absolute magnitudes. No systematic difference can be seen between LRDs and LBDs, and they occupy very similar locus in the EW(Ly)- plot as normal SFGs at the same radshift range identified with VLT/MUSE (J. Kerutt et al., 2022). Considering AGN-powered Ly emission can be more luminous than normal SFGs (e.g., D. Sobral et al., 2018), there seems no strong evidence supporting that the (narrow) Ly lines in LRDs or LBDs are mainly powered by AGN accretion. We should caution that, however, our search for the Ly emission with NIRSpec/Prism spectra could be incomplete particularly for low EW(Ly) sources due to the low spectral resolution of the Prism at these short wavelengths. It is also difficult to compare the intrinsic emission line profiles between Ly and H narrow component due to the low spectral resolution, which could otherwise be informative to probe the origin of the Ly line engine and escape mechanisms (S. Baek & A. Ferrara, 2013; Y. Yang et al., 2014). Further large spectroscopic campaigns of LRDs and LBDs at rest UV wavelengths with higher spectral resolution are needed to improve the statistics.
@@ IV.1 Constraint on LRD models from spectroscopic observations
In this section, we discuss the constraint on the models for LRDs from the observational results presented in the previous section. The observed correlations between the UV continuum and H lines, together with their luminosity ratios and Ly occurrence rates, provide key diagnostics of the dominant ionizing sources. We examine whether various envelope models (or “BH*” models) in the literature can consistently account for these observational constraints and assess their relative plausibility. We mainly focus on the envelope models proposed in literature (see Figure 1), but also briefly mention the possibility of other models than envelope models.
@@
(e.g., D. Kido et al., 2025; M. C. Begelman & J. Dexter, 2025).
@@
We also note that the stellar origin for the UV continuum can account for observations of LRDs spatially resolved in the rest-frame UV, comprising of the samples reported in the literature (e.g., P. Rinaldi et al., 2025; V. Kokorev et al., 2025). The remaining LRDs are spatially unresolved, including gravitationally lensed sources with tighter size limits () yet show V-shaped SEDs (e.g., L. J. Furtak et al., 2024b; C.-H. Chen et al., 2025), which suggests that the UV emission mainly originates from young compact stellar clusters in the nuclear region, as described in this model (Figure 1a). Overall, this model provides a self-consistent explanation for the observational results and is thus the most favored among the envelope models considered here.
@@
(Figure 1b; R. P. Naidu et al. 2025; A. de Graaff et al. 2025a).
This scenario can explain the tight correlation of the UV and the narrow H luminosity in LRDs. The stellar origin of the UV light in this model can also account for the morphology similar to the blackbody envelope model case.
However, it does not naturally account for the observed correlation between the UV continuum and the broad H luminosity.
@@
predominantly powered by the AGN accretion disk (e.g., X. Lin et al., 2025)111The non-spherical envelope model is defined as the case where both the UV continuum and the reprocessed emission lines powered by the AGN can escape through low-density, polar directions. Although X. Lin et al. (2025) adopt a non-spherical gas geometry and provide a schematic illustration, we here note that they do not explicitly specify the origin of the UV continuum, treating it as an open question..
@@
A possible explanation for the observed enhancement of H-to-UV luminosity ratio under this hypothesis is a higher gas covering fraction in the nuclear region during the LRD phase. Near the surface of the dense gaseous envelope (on spatial scales comparable to the BLR sizes inferred from reverberation mapping in nearby AGNs), BLR clouds occupy a large fraction of the solid angle of the radiation-escape funnel, thereby reprocessing ionizing photons into line emission more efficiently (e.g., R. Maiolino et al., 2025; Z. Yan et al., 2025). Therefore, the non-spherical envelope model could explain the different UV-to-H luminosity ratios between LRDs and LBDs if the BLRs during the LRD phase are assumed to have higher covering fractions and larger gas column densities than normal unobscured AGNs (LBDs). However, this assumption introduces a larger scatter in the UV-H relation and requires LBDs to have lower covering fractions. It also remains physically unclear why the offset between LRDs and LBDs appears at a factor of .
In any case, a non-spherical envelope phase may still play an important role during the evolutionary transition from LRDs to normal AGNs (see e.g., R. M. Mérida et al., 2025, for examples of transition-phase LRDs). The continuous distribution observed in the color–color diagram (Figure 2a) from LRDs to LBDs suggests that LRDs can gradually evolve into normal Type 1 AGNs (see also, e.g., G. Barro et al., 2025, for a similar discussion on the evolutionary stages of LRDs). Such a transition could naturally proceed through a non-spherical envelope phase, providing a physical link to the unified model of AGNs (e.g., H. Netzer, 2015).
@@
Although recent theoretical and observational works often focus on envelope models, they were originally proposed to explain a subset of LRDs with extreme properties such as deep Balmer breaks, Balmer absorption lines, exponential wing line profiles, or blackbody-like continuum shape from rest optical to infrared (e.g., R. P. Naidu et al., 2025; V. Rusakov et al., 2025a; X. Lin et al., 2025). It is thus possible that the dense gas envelope is needed only for these extreme LRDs and different physics may lie behind the overall normal population of LRDs.